Schedule of Cash Collections of Accounts Receivable

Furry Friends Supplies Inc., a pet wholesale supplier, was organized on May 1. Projected sales for each of the first three months of operations are as follows:

May $320,000
June 460,000
July 630,000

All sales are on account. 57 percent of sales are expected to be collected in the month of the sale, 32% in the month following the sale, and the remainder in the second month following the sale.

Prepare a schedule indicating cash collections from sales for May, June, and July.

Furry Friends Supplies Inc.
Schedule of Collections from Sales
For the Three Months Ending May 31
May June July
May sales on account:
Collected in May
Collected in June
Collected in July
June sales on account:
Collected in June
Collected in July
July sales on account:
Collected in July
Total cash collected $ $ $
Answer #1
Furry Friends Supplies Inc.
Schedule of Collections from Sales
For the Three Months Ending May 31
May June July
May sales on account:
Collected in May $1,82,400.00
Collected in June $1,02,400.00
Collected in July $   35,200.00
June sales on account:
Collected in June $2,62,200.00
Collected in July $1,47,200.00
July sales on account:
Collected in July $3,59,100.00
Total cash collected $1,82,400.00 $3,64,600.00 $5,41,500.00

Working

Furry Friends Supplies Inc.
Schedule of Collections from Sales
For the Three Months Ending May 31
May June July
May sales on account:
Collected in May =320000*57%
Collected in June =320000*32%
Collected in July =320000*11%
June sales on account:
Collected in June =460000*57%
Collected in July =460000*32%
July sales on account:
Collected in July =630000*57%